# Federated-inspired Single-cell Batch Integration in Latent Space

### v1

@@ 3.3 Local Objective with Proximal Regularization
$\mathcal{L}_{b}$ $=\mathbb{E}_{z\sim Z_{b}}\left[\|\gamma_{b}\odot z+\beta_{b}-z\|_{2}^{2}\right]$
$+\mu\left(\|\gamma_{b}-\gamma_{b}^{(t)}\|_{2}^{2}+\|\beta_{b}-\beta_{b}^{(t)}\|_{2}^{2}\right)$
@@ 4.2.1 Comparison with Baselines
Data Method Batch correction Bio. conservation Aggregate Added Time Two-batch PBMC PCA 0.3262 0.6601 0.5266 PCA-scBatchProx 0.3266 0.6603 0.5268 6.05 ICA 0.5315 0.6706 0.6150 ICA-scBatchProx 0.5328 0.6703 0.6153 6.13 FA 0.5095 0.6558 0.5973 FA-scBatchProx 0.5103 0.6788 0.6114 3.05 scVI 0.8047 0.7239 0.7562 scVI-scBatchProx 0.8280 0.7246 0.7660 5.05 LDVAE 0.6150 0.7331 0.6859 LDVAE-scBatchProx 0.6475 0.7305 0.6973 2.63 scANVI 0.8269 0.7359 0.7723 scANVI-scBatchProx 0.8426 0.7344 0.7777 2.94 HPMS PCA 0.3087 0.6473 0.5119 PCA-scBatchProx 0.3087 0.6522 0.5148 3.33 ICA 0.3433 0.6056 0.5007 ICA-scBatchProx 0.3440 0.6411 0.5223 2.93 FA 0.3074 0.6168 0.4930 FA-scBatchProx 0.3074 0.6169 0.4931 3.75 scVI 0.5447 0.6955 0.6352 scVI-scBatchProx 0.5456 0.6965 0.6361 3.59 LDVAE 0.5027 0.7091 0.6265 LDVAE-scBatchProx 0.5069 0.7095 0.6285 2.59 scANVI 0.5518 0.7320 0.6599 scANVI-scBatchProx 0.5566 0.7329 0.6624 3.92
@@ B.3 Aggregate Scores
Dataset Method Batch correction Bio. conservation Aggregate Two-batch PBMC PCA 0.3262 0.6601 0.5266 PCA-scBatchProx 0.3481 0.6515 0.5301 ICA 0.5315 0.6706 0.6150 ICA-scBatchProx 0.5455 0.6198 0.5900 FA 0.5095 0.6558 0.5973 FA-scBatchProx 0.5109 0.6803 0.6126 scVI 0.8047 0.7239 0.7562 scVI-scBatchProx 0.8221 0.7247 0.7636 LDVAE 0.6150 0.7331 0.6859 LDVAE-scBatchProx 0.6484 0.7303 0.6975 scANVI 0.8269 0.7359 0.7723 scANVI-scBatchProx 0.8066 0.7536 0.7748 HPMS PCA 0.3087 0.6473 0.5119 PCA-scBatchProx 0.3087 0.6522 0.5148 ICA 0.3433 0.6056 0.5007 ICA-scBatchProx 0.3740 0.6055 0.5129 FA 0.3074 0.6168 0.4930 FA-scBatchProx 0.3090 0.5447 0.4504 scVI 0.5447 0.6955 0.6352 scVI-scBatchProx 0.5445 0.6977 0.6364 LDVAE 0.5027 0.7091 0.6265 LDVAE-scBatchProx 0.5112 0.7421 0.6497 scANVI 0.5518 0.7320 0.6599 scANVI-scBatchProx 0.5504 0.7274 0.6566

### v2

# arXiv:2602.00423v2[cs.LG] 27 Feb 2026

@@ Abstract
1Department of Computer Science and Software Engineering, Auburn University
2Harrison College of Pharmacy, Auburn University
3College of Forestry, Wildlife and Environment, Auburn University
*Co-corresponding authors
hqn0001@auburn.edu   zzy0065@auburn.edu   hzc0134@auburn.edu   wzk0004@auburn.edu   jqwang@auburn.edu
@@ 3.3 Local Objective with Proximal Regularization
$\mathcal{L}_{b}$ $=\mathbb{E}_{z\sim Z_{b}}\left[\|\gamma_{b}\odot z+\beta_{b}-z\|_{2}^{2}\right]+\mu\left(\|\gamma_{b}-\gamma_{b}^{(t)}\|_{2}^{2}+\|\beta_{b}-\beta_{b}^{(t)}\|_{2}^{2}\right)$
@@ 4.2.1 Comparison with Baselines
Data Method Batch correction Bio. conservation Aggregate Added Time
Two-batch PBMC PCA 0.3262 0.6601 0.5266
PCA-scBatchProx 0.3266 0.6603 0.5268 6.05
ICA 0.5315 0.6706 0.6150
ICA-scBatchProx 0.5328 0.6703 0.6153 6.13
FA 0.5095 0.6558 0.5973
FA-scBatchProx 0.5103 0.6788 0.6114 3.05
scVI 0.8047 0.7239 0.7562
scVI-scBatchProx 0.8280 0.7246 0.7660 5.05
LDVAE 0.6150 0.7331 0.6859
LDVAE-scBatchProx 0.6475 0.7305 0.6973 2.63
scANVI 0.8269 0.7359 0.7723
scANVI-scBatchProx 0.8426 0.7344 0.7777 2.94
HPMS PCA 0.3087 0.6473 0.5119
PCA-scBatchProx 0.3087 0.6522 0.5148 3.33
ICA 0.3433 0.6056 0.5007
ICA-scBatchProx 0.3440 0.6411 0.5223 2.93
FA 0.3074 0.6168 0.4930
FA-scBatchProx 0.3074 0.6169 0.4931 3.75
scVI 0.5447 0.6955 0.6352
scVI-scBatchProx 0.5456 0.6965 0.6361 3.59
LDVAE 0.5027 0.7091 0.6265
LDVAE-scBatchProx 0.5069 0.7095 0.6285 2.59
scANVI 0.5518 0.7320 0.6599
scANVI-scBatchProx 0.5566 0.7329 0.6624 3.92
@@ B.3 Aggregate Scores
Dataset Method Batch correction Bio. conservation Aggregate
Two-batch PBMC PCA 0.3262 0.6601 0.5266
PCA-scBatchProx 0.3481 0.6515 0.5301
ICA 0.5315 0.6706 0.6150
ICA-scBatchProx 0.5455 0.6198 0.5900
FA 0.5095 0.6558 0.5973
FA-scBatchProx 0.5109 0.6803 0.6126
scVI 0.8047 0.7239 0.7562
scVI-scBatchProx 0.8221 0.7247 0.7636
LDVAE 0.6150 0.7331 0.6859
LDVAE-scBatchProx 0.6484 0.7303 0.6975
scANVI 0.8269 0.7359 0.7723
scANVI-scBatchProx 0.8066 0.7536 0.7748
HPMS PCA 0.3087 0.6473 0.5119
PCA-scBatchProx 0.3087 0.6522 0.5148
ICA 0.3433 0.6056 0.5007
ICA-scBatchProx 0.3740 0.6055 0.5129
FA 0.3074 0.6168 0.4930
FA-scBatchProx 0.3090 0.5447 0.4504
scVI 0.5447 0.6955 0.6352
scVI-scBatchProx 0.5445 0.6977 0.6364
LDVAE 0.5027 0.7091 0.6265
LDVAE-scBatchProx 0.5112 0.7421 0.6497
scANVI 0.5518 0.7320 0.6599
scANVI-scBatchProx 0.5504 0.7274 0.6566